# Generating Completions for Fragmented Broca’s Aphasic Sentences Using Large Language Models

### v1

@@ 4 Discussion
In this study, we aimed to explore the potential of reconstructing fragmented Broca’s aphasic language into full utterances using LLMs fine-tuned on synthetic Broca’s aphasic data. Using a rule-based system, we generated synthetic data based on transcribed spoken language from neurotypical individuals. Our synthetic data mirrors the characteristics of Broca’s aphasic speech. Subsequently, we fine-tuned four LLMs on this synthetic data to evaluate their ability to complete fragmented aphasic sentences. To the best of our knowledge, this is the first study to evaluate LLMs on the task of reconstructing authentic Broca’s aphasic sentences.
@@
4.1 Automatic fragmented sentence completion
@@ 4.1.1 Synthetic data
We found that the models performed well on synthetic aphasic data, generating coherent and contextually appropriate completions. The T5-base and Flan-T5-XL models showed consistent performance in the synthetic setting, with Flan-T5-XL exhibiting a slightly better use of context and producing slightly better completions.
@@ 4.1.2 Authentic data
When tested on authentic Broca’s aphasic sentences, the models’ performance varied. The models were able to generate reasonable completions based on visual inspections, with better results observed for longer input sentences, which provide more contextual information, resulting in improved model performance. Visual inspections indicated that Flan-T5-XL generated more coherent and contextually relevant completions compared to T5-base, particularly for longer utterances. These findings suggest that sequence-to-sequence LLMs are effective in generating meaningful and contextually appropriate completions for fragmented Broca’s aphasic sentences, indicating their potential for practical applications in language rehabilitation.
@@ 4.1.3 Implications
The demonstrated ability of sequence-to-sequence LLMs to generate reasonable completions for fragmented Broca’s aphasic sentences highlights their potential in advancing communicative technologies for non-fluent aphasic patients. These models could become valuable tools in speech and language therapy by providing contextually appropriate sentence completions, potentially helping individuals with Broca’s aphasia convey their thoughts more effectively.
@@ 4.2 Limitations
The computational resources required for training and fine-tuning large transformer-based models are substantial. This limitation may restrict the accessibility and scalability of the proposed approaches, specifically for researchers and practitioners with limited computational resources.
Finally, recovering the intended meaning of fragmented aphasic sentences without extensive context or direct interaction with the individual is inherently challenging. For instance, the fragment “He book table” could be interpreted in multiple ways such as “He put the book on the table”, “He took the book from the table” or “He wants the book on the table.” This variability emphasises the challenge that a sentence cannot always be accurately recovered without additional grounding such as understanding the broader conversation or using external data such as visual cues.
@@ 4.3 Future work
In terms of synthetic Broca’s aphasic data generation, one route for future research is the exploration of data augmentation techniques tailored to fragmented aphasic sentences, improving generalisation capabilities of the models on real-world Broca’s aphasic speech.
Another interesting direction for future work is studying the domain influence, possibly enabling the creation of task-specific models based on the patient’s interests or hobbies. The incorporation of domain knowledge has demonstrated improved model performance in NLP, including machine translation [37] and NLP-based educational applications [38].
Finally, future research could ask native English speakers to produce corresponding target completions for the authentic Broca’s aphasic utterances from AphasiaBank. The availability of such ground-truth data is crucial for a quantitative evaluation of models’ completions of the authentic Broca’s aphasic sentences. The creation of a reliable benchmark would provide a reference for evaluating how well the sentence completion models handle typical errors and omissions in Broca’s aphasia. Moving beyond the sentence completion task, such data would contribute immensely to creating a broader benchmark for related tasks in the field of aphasiology.
@@ 5 Conclusion
In summary, we conclude that sequence-to-sequence large language models exhibit the capability of reconstructing fragmented sentences produced by individuals with Broca’s aphasia. The models generate reasonable and contextually appropriate completions, denoting their potential as valuable tools in speech and language therapy. These findings support the ongoing development and refinement of machine learning models tailored to assist in communication for those affected by language impairments. The broader potential to create synthetic data for various clinical populations opens avenues for improving their communicative independence and quality of life.

### v2

@@ 4 Discussion
In this study, we aimed to explore the potential of reconstructing Broca’s aphasic language into full utterances using LLMs fine-tuned on synthetic Broca’s aphasic data. Using a rule-based system, we generated synthetic data based on transcribed spoken language from neurotypical individuals. Our synthetic data mirrors the characteristics of Broca’s aphasic speech. Subsequently, we fine-tuned four LLMs on this synthetic data to evaluate their ability to complete agrammatic aphasic sentences. To the best of our knowledge, this is the first study to evaluate LLMs on the task of reconstructing authentic Broca’s aphasic sentences.
@@
4.1 Automatic agrammatic sentence completion
@@ 4.1.1 Synthetic data
We found that the models performed well on synthetic aphasic data, generating coherent and contextually appropriate completions. The T5-base and Flan-T5-XL models showed consistent performance in the synthetic setting, with Flan-T5-XL exhibiting a slightly better use of context and producing slightly better completions. These findings are consistent with the study of[30], who demonstrated the utility of synthetic data for training LLMs to reconstruct agrammatic aphasic sentences. However, while [30] used written text from the C4 corpus to generate synthetic aphasic utterances, our approach is based on spoken language, which better reflects conversational contexts encountered by PWA. Additionally, unlike [30] who evaluated their model solely on synthetic data, we extended the evaluation to authentic aphasic sentences, addressing a key limitation in their study. This broader evaluation provides a more realistic assessment of the models’ applicability to clinical scenarios.
@@ 4.1.2 Authentic data
When tested on authentic Broca’s aphasic sentences, the models’ performance varied. The models were able to generate reasonable completions based on visual inspections, with better results observed for longer input sentences, which provide more contextual information, resulting in improved model performance. Visual inspections indicated that Flan-T5-XL generated more coherent and contextually relevant completions compared to T5-base, particularly for longer utterances. In contrast to [30] study, which did not test their approach on real aphasic data, our findings demonstrate that fine-tuned LLMs can generalize effectively to authentic agrammatic sentences produced by individuals with Broca’s aphasia. This step highlights the potential and the practical relevance of such models for rehabilitation purposes. While previous AI applications in aphasia have largely focused on diagnostic tasks or static assistive technologies [27, 28, 29, 3], our results suggest that LLM-based systems have the potential to provide adaptive solutions for reconstructing agrammatic speech into complete sentences. Incorporating longer input utterances further improves model performance, suggesting the importance of designing systems capable of leveraging contextual information to support real-world communication needs for PWA and their caregivers.
@@ 4.1.3 Implications
The demonstrated ability of sequence-to-sequence LLMs to generate reasonable completions for Broca’s aphasic sentences highlights their potential in advancing communicative technologies for non-fluent aphasic patients. These models could become valuable tools in speech and language therapy by providing contextually appropriate sentence completions, potentially helping individuals with Broca’s aphasia convey their thoughts more effectively.
@@ 4.2 Limitations
The computational resources required for training and fine-tuning large transformer-based models are substantial. This limitation may restrict the accessibility and scalability of the proposed approaches, specifically for researchers and practitioners with limited computational resources. Moreover, we only trained and tested the models on one high-resource language, English, and future work should determine whether our proposed method can also be used in other languages.
Finally, recovering the intended meaning of agrammatic aphasic sentences without extensive context or direct interaction with the individual is inherently challenging. For instance, the fragment “He book table” could be interpreted in multiple ways such as “He put the book on the table”, “He took the book from the table” or “He wants the book on the table.” This variability emphasises the challenge that a sentence cannot always be accurately recovered without additional grounding such as understanding the broader conversation or using external data such as visual cues.
@@ 4.3 Future work
In terms of synthetic Broca’s aphasic data generation, one route for future research is the exploration of data augmentation techniques tailored to agrammatic aphasic sentences, improving generalisation capabilities of the models on real-world Broca’s aphasic speech.
Another interesting direction for future work is studying the domain influence, possibly enabling the creation of task-specific models based on the patient’s interests or hobbies. The incorporation of domain knowledge has demonstrated improved model performance in NLP, including machine translation [53] and NLP-based educational applications [54]. In terms of training setup, it could be promising to explore in-context learning methods with the same models, and compare the resulting scores with those reported in our study (similar, to, e.g., [55]). Our preliminary experiments with FLAN-T5-XL in a ten-shot learning setup suggest that fine-tuning methods yield better performance (on ChrF and Rouge), but experimenting with a variety of prompts may be necessary in the future.
Finally, we recognize that the observed correct sentence completion rate is not sufficient for deployment as a ready-to-use communication aid in clinical settings. Nonetheless, this work represents an initial step toward building such applications, since there are currently no existing benchmarks for performance on authentic aphasic sentences. Therefore, future research could ask native English speakers to produce corresponding target completions for the authentic Broca’s aphasic utterances from AphasiaBank. The availability of such ground-truth data is crucial for a quantitative evaluation of models’ completions of the authentic Broca’s aphasic sentences. The creation of a reliable benchmark would provide a reference for evaluating how well the sentence completion models handle typical errors and omissions in Broca’s aphasia. Moving beyond the sentence completion task, such data would contribute immensely to creating a broader benchmark for related tasks in the field of aphasiology.
@@ 5 Conclusion
In summary, we conclude that sequence-to-sequence large language models exhibit the capability of reconstructing agrammatic sentences produced by individuals with Broca’s aphasia. The models generate reasonable and contextually appropriate completions, denoting their potential as valuable tools in speech and language therapy. These findings support the ongoing development and refinement of machine learning models tailored to assist in communication for those affected by language impairments. The broader potential to create synthetic data for various clinical populations opens avenues for improving their communicative independence and quality of life.